A WHOLE NUMBER IS DIVISIBLE BY 3, IF IT'S SUM IS DIVISIBLE BY 3..NOT SURE HOW TO DO THIS PLEASE HELP..
Answer by Cintchr(475)

This rule works for dividing by nine as well.
take the number 345. Add the digits up.
3 + 4 + 5 = 12
look at the answer ... 12. 12 is divisible by 3, so the number 345 is as well.
take a HUGE number .... 34,826,406,723,467
add the digits ...
3+4+8+2+6+4+0+6+7+2+3+4+6+7=62
we can now take 62 ... 6 + 2 = 8 .
3 can not go into 8 ... so it can not go into 34,826,406,723,467.
The same holds true with the number 9. but the sum of the numbers has to be divisible by 9.
85,716
8+5+7+1+6=27
9 goes into 27, so it goes into 85,716